OpenStudy (anonymous):

−8 −5x + 2 > −38 im lost on certain parts help plz

OpenStudy (anonymous):

you solve it like an equation. Get the x on one side and everything else on the other side. What are you confused about ?

OpenStudy (anonymous):

btw in between -8 and -8 is a greater and equal too sign

OpenStudy (anonymous):

like do start by addin -2 on both sides?

OpenStudy (anonymous):

wait....is the problem -8 >= -5x + 2 > -38 ?

OpenStudy (anonymous):

no its -8 great equal to -5x +2 >-28

OpenStudy (anonymous):

-8 >= -5x + 2 >-28 ....is that the correct problem ?

OpenStudy (anonymous):

-8≥-5x+2>-38

OpenStudy (anonymous):

i added -2 to both sides first

OpenStudy (anonymous):

which gave me -10≥-5x-40

OpenStudy (anonymous):

-10≥-5x>-40

OpenStudy (anonymous):

when i divide just to make sure both signs are guna flip right?

OpenStudy (anonymous):

ok....your dealing with a compound inequality. You have to take it apartand make two problems. -8 >= 5x + 2 -8 - 2 >= 5x -10 >= 5x -10/5 <= x -2 <= x now the second part... -5x + 2 > - 38 -5x > - 38 - 2 -5x > -40 x < -40/-5 x < 8 so your answer is -2 <=x, and x < 8 or...-2 <= x < 8

OpenStudy (anonymous):

im trying it right now

OpenStudy (anonymous):

wait...I made a mistake... the first part.... -8 >= 5x + 2 -8 -2 >= 5x -10 >= 5x -10/5 >= x -2 >= x so your answer would be... -2 >=x, and x < 8

OpenStudy (anonymous):

and for the if it was on a number line it would be like -2 open circle to the right and open circle to the left of 8 right?

OpenStudy (anonymous):

closed circle on -2, shaded to the left. Open circle on 8, and shaded to the left.

OpenStudy (anonymous):

my bad....closed circle on -2 and open circle on 8, shading in between. (I think)

OpenStudy (anonymous):

it is i just checked

OpenStudy (anonymous):

I get confused with the negatives.....it is a closed circle on -2 because there is an equal sign involved. x is less then -2. So, I am not sure if the shading is to the right or left ? Open circle on 8 and shaded to the left

OpenStudy (anonymous):

you understand how you have to make two problems out of a compound inequality ?

OpenStudy (anonymous):

yup thanks to your help and a review on the lesson its all clear now thanks :)

OpenStudy (anonymous):

your welcome :) I just hope I didn't confuse you more with all of my " mess ups "

OpenStudy (anonymous):

haha its ok :)